Government-wide statements report on assets and liabilities that are denied recognition on funds statements. Entrepreneurs Consultants, a state agency, was estab- lished to provide consulting services to small businesses. It maintains only a single general fund and accounts for its activities on a modified accrual basis. During its first month of operations, the association engaged in, or was affected by, the following transactions and events: 1. It received an unassigned grant of $100,000. 2. It purchased five computers at $2,000 each. 3. It paid wages and salaries of $6,000. 4. Itborrowed$24,000fromabanktoenableittopurchase an automobile. It gave the bank a long-term note. 5. It purchased the automobile for $24,000. 6. It made its first payment on the note—interest of $200. 7. It destroyed one of its computers in an accident. The computer was not insured. a. Prepare journal entries in the general fund to record each of the transactions or other events. b. Prepare a balance sheet and a statement of revenues and expenditures for the general fund. c. Prepareagovernment-widestatementofnetposition (balance sheet) and statement of activities. These should be on a full accrual basis. Assume that the capital assets have a useful life of five years and that no

depreciation is to be charged on the computer that was destroyed.    THIS QUESTION IS SOLVED ALREADY BUT I WOULD LIKE TO RECONCILE THE CASH $ 83,800 AND THE FUND BALANCE $ 85,400. THE ANSWER IS ALREADY ON CHEGG SO YOU CAN TAKE A LOOK AND JUST RECONCILE THE TWO AMOUNTS.
Answer #1

1. Calculation of Cash balance

Particulars Amount
Grant Received 100,000
Less: Computers Purchased (10,000)
Less: Wages & Salaries (6,000)
Add: Loan Received 24,000
Less: Automobile Purchased (24,000)
Less: Interest paid on loan (200)
Cash Balance 83,800

1. Calculation of Fund balance

Particulars Amount
Grant Received 100,000
Less: Wages & Salaries Paid (6,000)
Add: Bank Loan 24,000
Less: Automobile Purchase (24,000)
Less: Interest on note (200)
Less: Net Value of Computers (NOTE) (8,400)
Fund Balance 85,400

NOTE: Calculation of net value of assets

Computer Purchased = $10,000

Depreciation on computers = (4 x 2,000) / 5 years = $1,600

Net Value of computer = 10,000 - 1,600 = $8,400